OpenStudy (anonymous):

One leg of a right triangle is 10m less than the hypotenuse, and the other leg is 5m less than the hypotenuse. find the length of all three sides

OpenStudy (hitaro9):

Okay. It always helps to set up an equation. You should know that a^2 + b^2 =c^2, yes?

OpenStudy (anonymous):

yeah i dd that

OpenStudy (hitaro9):

If c is the hypotenuse, and a and b are the two side. Then the a is going to be c-10 (a = c-10m) and b is going to be c-5.

OpenStudy (hitaro9):

Understand that?

OpenStudy (anonymous):

yeah

OpenStudy (hitaro9):

Alright. From there its just a whole bunch of multiplication after you plug it back in to the equation. a= c-10 b = c-5 a^2 + b^2 = c^2 (c-10)^2 + (c-5)^2 = c^2

OpenStudy (hitaro9):

You'll want to do two FOILS, and then do some adding/ subtracting from both sides to get c. From there you'll know a = c-10 and b = c-5, so those will both be really simple to do.

OpenStudy (hitaro9):

Hopefully that all makes sense. All you've got left to do now is a little tedious multiplication and algebra. Any questions?

OpenStudy (anonymous):

No, it makes sense thanks.

OpenStudy (hitaro9):

Alright. Good luck with the problem. That equation looks pretty ugly/ easy to make a careless mistake on.